A big title

me

September 24, 2003

A little header

me

September 24, 2003

BOLD Italic boxed typewriter

1. first item

2. second item

1 a section

text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text

new paragraph with math f(x) = 3x^{2} + \frac{sin(x)}{y_{1}^{2α}}.

∫_{a}^{b} f(x)dx = ∑_{i=3}^{2n−2} f̂(3x) = \{ \begin{matrix}3 & x = 2 \\ 0 & otherwise \end{matrix}

blah

[ \begin{matrix}1 & 2 & 3 \\ 4 & ϵ & 5 \end{matrix} ] [ \begin{matrix}1 & 2 & 3 \\ 4 & ϵ & 5 \\ 4 & ϵ & ( \begin{matrix}5 & 6 \\ 7 & 8 \end{matrix} ) \end{matrix} ]  (1)

I like referring to (1).

v_{3}(x, y) = {(d_{i}^{1}, x_{i}^{1}(x), y_{i}^{1}(y))}_{i=1}^{r_{1}} ∪ {(d_{j}^{2}, x_{j}^{2}(x), y_{j}^{2}(y))}_{j=1}^{r_{2}}  (2)
= {(d_{k}^{3}, x_{k}^{3}(x), y_{k}^{3}(y))}_{k=1}^{r_{3}} .  (3)